OpenStudy (anonymous):

2/ (x(x+2)). As lim of x approaches 0 ^- ,0^+ , and 0

OpenStudy (anonymous):

Is it -infinity ,+infinity and ,doesnt exist respectively ?

OpenStudy (mimi_x3):

\[\lim_{x \rightarrow 0} \frac{2}{x(x+2)} => \lim_{x \rightarrow 0} \frac{2}{x^2+2x} => \lim_{x \rightarrow 0} \frac{2/x^2}{x^2/x^2+2x/x^2}\] \[\lim_{x \rightarrow 0} \frac{2/x^2}{1+2/x} =0\]

OpenStudy (mimi_x3):

I mean = infinity

OpenStudy (mimi_x3):

not 0 lol

OpenStudy (anonymous):

Ah

OpenStudy (mimi_x3):

limit doesnt exist because the left hand limit does not equal the right hand limit.

OpenStudy (anonymous):

But what if x approaches 0 minus

OpenStudy (mimi_x3):

negative infinity.

OpenStudy (anonymous):

Now it makes sense, thanks

OpenStudy (mimi_x3):

No problem.